# BARE: Combining Base and Instruction-Tuned Language Models for Better Synthetic Data Generation

### v2

@@ 3.3 Investigation Results
Figure 3: Distribution of pairwise embedding cosine similarities for Llama-3.1-70B-Instruct and Llama-3.1-70B-Base generations for Enron spam emails. Base model distribution have more density in the low-similarity region and less density in the high-similarity region, indicating greater diversity. Note that frequency is of the cosine similarity score.
Figure 4: Topic coverage of Llama-3.1-70B-Instruct and Llama-3.1-70B-Base generations for Newsgroups. Coverage after $n=500$ generations. The base model covers all topics except one whereas the instruct model covers only $60\%$ of the topics.
@@
Looking at the pairwise cosine similarity distributions of the embeddings of the generated dataset in Figure 3 (and recalling Figure 2), the base distribution is consistently shifted to the left, indicating that the base model generations are more diverse and correspond with a decrease in average similarity (Table 1). The higher diversity of base models is more strikingly shown in the coverage of generated classes for Newsgroups (Figure 4), where over $500$ generations the Instruct model covers only 14 of the 20 classes. In contrast, the Base model covers every class except 1 (talk.religion.misc).
@@
Dataset Instruct Base
Enron 0.450 0.350
Newsgroups 0.256 0.162
GSM8K 0.421 0.313
LCB TOP 0.389 0.468
This trend is reflected in Table 1, with all domains except one showing results for base models with higher diversity. Upon inspection, we attribute the reversal in trend in LCB TOP to the base model repeating phrases from the examples in the prompt. This is related to potential issues with the quality of base model generations, which we discuss below.
@@
Method Enron GSM8K
GPT-4o Independent 0.574 0.427
GPT-4o Persona 0.580 0.308
GPT-4o Sequential 0.511 0.398
GPT-4o In-One 0.363 0.347
GPT-4o Dynamic Fewshot 0.511 0.463
Llama-3.1-70B-Base 0.350 0.313
Sampling Methods
We further explore data generation using several different sampling methods. As base models struggle to follow complex instructions, only instruct models are used for methods beyond independent sampling.
  • Independent (temperature) Sampling: The model generates $n$ samples to form a dataset.
  • Persona Prompting: The model is instructed to respond as a predefined persona (Fröhling et al., 2024).
  • Sequential Prompting: The model iteratively generates outputs distinct from previous ones.
  • In-One Prompting: The model is prompted to generate $k$ different entries in a single response (Zhang et al., 2024b).
  • Dynamic Few-Shot Examples: Few-shot examples are randomly selected for each call (breaking the assumption of low-data conditioned generation) (Li et al., 2022).
Table 2 shows that base models generally yield higher diversity than almost all prompting methods on the average embedding distance metric in two domains. One exception is persona prompting on GSM8K — though this diversity arises more from flavor text differences due to personas rather than actual content.
We thus find that base models are generally more diverse than instruct-tuned models and that temperature increases and prompting methods are insufficient to bridge the gap, motivating our usage of base models in the first stage of BARE.
Figure 5: Indistinguishability Rate for Llama-3.1-70B-Base and Instruct model data generation methods across various datasets. Llama-3.1-70B-Instruct is almost uniformly better at appearing high-quality alongside real-world data.
@@
Figure 5 presents our results. In general, the instruct-tuned model has a higher IR, indicating that it is better at producing generations that resemble high-quality data. Note that some IRs are well above 75%. This outcome is not unexpected: if a model consistently generates data that aligns with the most common patterns in the real-world distribution, it becomes difficult to distinguish from actual data. Since the real-world entries often also include non-modal (less frequent) samples, a discriminator tasked with identifying lower quality data may instead misclassify these less common real-world samples as synthetic.
As mentioned above, on LCB TOP the base model repeats phrases from examples in the prompt. This leads to a higher IR as the generations copy real-world examples. Thus, while individual base generations are technically more realistic, their shortcomings are captured in the diversity metric.
We therefore find that the superior instruction follow-up capabilities of instruct-tuned models can help generate more realistic data, motivating our usage of instruct-tuned models in the second stage of BARE.
@@ 4 Base-Refine (BARE)
Figure 6: Instruct-tuned models provide high-quality but low-diversity data, while base models provide low-quality but high-diversity data. BARE independently generates a diverse initial set of data points with a base model and refines each entry individually with an instruct-tuned model to create a high-quality, high-diversity dataset. In this example of a real grade school math problem generation, the Llama-3.1-70B-Base model hallucinates in its answer to its own question. The refiner (Llama-3.1-70B-Instruct) recognizes this and disambiguates the question and corrects the reasoning.
We leverage our insights to propose BARE, a practical synthetic data generation method combining the diversity of base models with the quality of instruct models. BARE uses a base model to generate an initial set of diverse but potentially lower quality data, after which an instruct-tuned model individually refines each example from the initial set – as shown in the example in Figure 6.
In the base step, the base model uses a prompt with minimal few-shot examples to generate. In the refine step, the instruct-tuned model is instructed to retain the concept of the base model generation while improving it according to specific criteria (e.g., realism, correctness). This retains the overall diversity of the base model generated set while exerting greater control over the quality of the final generation.
Importantly, few-shot examples are only required in the base step to ensure formatting (though they can also be included in the refine step), and only in small amounts. Thus, BARE is especially useful in domains with very limited real-world data as little human effort is required to generate a diverse high-quality synthetic dataset. In our experiments, we limit ourselves to just three few-shot examples. In addition, we intentionally use very general prompts for BARE to demonstrate its flexibility, underscoring the potential for even greater improvement with tailored prompts. Representative prompts are included in Appendix B.
@@ 5 Evaluation
We evaluate BARE for diversity, data quality, and downstream utility on the same domains and against the same baseline methods presented in Section 3. We implement BARE with Llama-3.1-70B-Base in the base generation stage and Llama-3.1-70B-Instruct in the refinement stage, and perform additional experiments with the Llama 3.1 8B family and with Llama 3.1 8/70B as the base model and GPT-4o as the refiner (Dubey et al., 2024; OpenAI, 2024c).
We follow generally the same sampling temperatures as in Section 3. However, for Enron and Newsgroups, we instead sample from Llama-3.1-8B-Instruct at temperature $0.7$ in order to maintain coherence in generations. We always sample the refine model at a temperature of $0.7$.
For evaluation of downstream utility for the generative tasks, we LoRA (Hu et al., 2021) fine-tune Llama-3.1-8B-Instruct for 4 epochs using the generated data, except for GSM8K where Llama-3.2-1B-Instruct (Dubey et al., 2024) is fine-tuned instead due to high baseline performance of the 8B model. Other fine-tuning hyperparameters are in Section A.1.1. The fine-tuned models are evaluated on a static $n=100$ test set for HotpotQA and PubMedQA, a static $n=500$ test set for GSM8K, and the full $n=442$ test set for LCB TOP. In this section, we will focus only on the generative tasks; the downstream evaluation process for classification tasks is presented in Section A.1.2 and detailed results for all domains can be found in Section A.2.
Figure 7: Distribution of pairwise embedding similarities for Llama-3.1-70B-Base and Llama 3.1 70B BARE generations for GSM8K (top) and LCB TOP (bottom). The Base and BARE distributions are extremely similar for both tasks, indicating that refinement retains the diversity of the base generations. Note that frequency is of the cosine similarity score.
@@
We begin by looking at the quality/diversity trends of BARE when compared with previous methods. From the histograms plotted in Figure 7, we can see that BARE effectively does not change the similarity distribution of generated data when compared to the base model at all – it is able to successfully retain the diversity of base generations. Detailed results with average embedding similarity scores can be found in Section A.2.
Figure 8: Indistinguishability Rate for Llama-3.1-70B-Base and BARE data generation across various datasets. BARE consistently improves the quality of data generated from base models.
Simultaneously though, looking at Figure 8, we can see that while retaining diversity, BARE leads to a monotonic increase in the IR for every domain - suggesting that it is able to lift the quality of the generations to be on par or in some cases even surpass directly sampling from an instruct model. Combined, this indicates that BARE is capable of leveraging the diversity of base models and quality of instruct-tuned models in its end generations.
Figure 9: Fine-tuned model accuracy results on HotpotQA, PubMedQA, GSM8K, and LCB TOP for prompting methods using GPT-4o. Prompting methods have mixed effects on downstream performance when compared against standard sampling, while using GPT-4o with simply the Llama-3.1-8B-Base model in BARE improves consistently over all prompting methods.
Figure 10: Accuracy of a Llama-3.1-8B-Instruct Model (HotpotQA, PubMedQA, LCB TOP) and Llama-3.2-1B-Instruct Model (GSM8K) finetuned on synthetic data generated using Base, Instruct, and BARE methods. Note that GSM8K 8B Instruct results are not shown as data generations were derailed.
@@
We now show the utility of BARE datasets as a whole. In Figure 10, we demonstrate the accuracy of a model fine-tuned on the datasets generated using different methods. Almost uniformly, BARE leads to greater improvements in downstream model quality than generating from either Base or Instruct models, and across all domains BARE-based data leads to the highest fine-tuned model accuracy.
Impressively, in Figure 9, beyond just base and instruct sampling, we show that enhancing GPT-4o with BARE using only a small base model like Llama-3.1-8B-Base produces monotonically better fine-tuning data than all prompting methods discussed in Section 3.3 that use only GPT-4o. These clear results showcase BARE’s ability to out-perform the existing methods for high-quality, diverse data generation.
Focusing on the RAFT domains, BARE improves upon the standard SOTA pipeline for fine-tuning LLMs for RAG by up to 18.4%, as seen with the 8B family on HotpotQA (standard RAFT is implemented in our Instruct generation results). BARE with both model families also outperforms existing RAFT pipelines on PubMedQA. While BARE with the Llama 3.1 70B family does not improve upon RAFT for HotpotQA, switching out the Llama-3.1-70B-Instruct refiner for GPT-4o does lead to an improvement (Section A.2).
On GSM8K, BARE is the only method that provides useful training data. The un-trained model performance was 21.8%, and fine-tuning on BARE generated data achieves accuracies of 24.9% and 32.8% with the Llama 8B and 70B families, respectively. Accuracy when training with data generated by single model methods either decreased accuracy or had little difference. In fact, training on data generated by Llama-3.1-70B-Instruct led to an accuracy of just 17.8%, which BARE with Llama-3.1-70B-Base refined by GPT-4o outperforms by 101% (35.8% accuracy).
@@
Similar to HotpotQA, on LCB TOP, BARE with the Llama 3.1 70B family shows less of an improvement than the 70B-instruct-only data generation method. Again, using GPT-4o as a refiner instead of Llama-3.1-70B-Instruct gives stronger results. This suggests that with the right refiner choice, BARE consistently provides gains. We show detailed results of refining with GPT-4o in all domains in Section A.2, emphasizing BARE’s consistently strong data generation capabilities.
We also find that temperature ablations do not meaningfully affect the utility of instruct-tuned model generations, especially when compared to gains using BARE. For details, see Section A.3.
@@
A possible challenge is that the gains in performance are due to a multi-step pipeline rather than combining a base and instruct-tuned model. To demonstrate the importance of using a base model in BARE, we performed an ablation where an instruct-tuned model’s generations are refined by an instruct-tuned model on the GSM8K task. The accuracy on the test set after switching Llama-3.1-70B-Base to Instruct in the first step drops from $29.8\%$ to $25.4\%$. This trend held when GPT-4o was the refiner as well, with a drop from $35.8\%$ to $30.8\%$. At the same time, the pattern of little change in the average pairwise similarity before and after refinement remains, indicating that diversity must be introduced in the first stage. Having earlier established that base models are most effective at diversity, we conclude that the use of base models is a necessary component of BARE. Detailed results from this ablation are in Section A.4.
@@ 6 Conclusion
6.1 Key Contributions
In summary, in this work we quantitatively investigated the quality and diversity of various synthetic data generation methods. Through this investigation, we find that base models are generally more diverse than instruct-tuned models while instruct-tuned models produce higher quality output than base models, validating hypotheses in the research community. These insights motivate the design of a better system to generate synthetic data, BARE.
Through extensive experiments, we validate the importance of each step in BARE and demonstrate its ability to preserve base model diversity while enhancing output quality. Moreover, by fine-tuning on BARE-generated data for various domains, we underscore BARE’s practical utility, consistently outperforming existing synthetic data generation methods on downstream tasks such as GSM8K and LiveCodeBench, in addition to RAFT, for which we set a new SOTA.
6.2 Future Work
There are a lot of exciting directions for future work. For one, BARE is not necessarily the only way to elicit diversity in the synthetic data generation process. We were able to achieve all of our results with essentially basic prompting, leading us to believe there is room for even more improvement if one was to, for example, fine-tune the refiner specifically for this use case or use stronger refiner models. Introducing additional stages to BARE could also lead to improvements (e.g., additional refinement steps). Beyond multi-stage systems, the design space for diversity is vast – in one direction, training instruct-tuned models with different objective functions that encourage entropy is another area of exploration.
Furthermore, the implications of a lack of high-quality, diverse generations from LLMs go beyond just synthetic data. For example, much of the work in inference time compute relies on models being able to generate sufficiently different possible trajectories and using these diverse trajectories (exploration) with a feedback signal to improve reasoning capabilities (Zhang et al., 2024a; Zelikman et al., 2022), yet only a few methods look explicitly at generator diversity (Wang et al., 2024).
Lastly, the same BARE method demonstrated in this paper for synthetic training data can also be used with no modifications to generate synthetic evaluation sets, which are especially valuable in so many real-world domains with low-data availability.
Acknowledgements
@@
Sky Computing Lab is supported by gifts from Accenture, AMD, Anyscale, Cisco, Google, IBM, Intel, Intesa Sanpaolo, Lambda, Microsoft, NVIDIA, Samsung SDS, SAP, and VMware.
@@
Appendix A Additional Results
A.1 Downstream Evaluation - Additional Details
A.1.1 Fine-Tuning Task Hyperparameters

### v3

@@ 3.3 Investigation Results
From the pairwise cosine similarity distributions of the embeddings in Figure 3 (and recalling Figure 2), we see the base distribution is consistently further to the left, indicating that the base model generations are more diverse. The better diversity of base models is further reflected in Table 1, with all domains except one showing lower mean similarity for base models, indicating higher diversity. Upon inspection, we attribute the reversal in trend in LCB TOP to the base model repeating phrases from the examples. This is related to potential issues with the quality of base model generations, which we discuss below.
@@
Enron Newsgroups GSM8K LCB TOP
Instruct 0.450 0.256 0.421 0.389
Base 0.350 0.162 0.313 0.468
In addition to repeated independent sampling from the instruction-tuned model, we further explore diversity-eliciting prompting for data generation. We do not explore such methods on base models due to their poor instruction-following capabilities. The methods we examine are:
  • Persona Prompting: Model responds as a predefined persona (Fröhling et al., 2024).
  • Sequential Prompting: Model iteratively generates outputs distinct from previous ones.
  • In-One Prompting: Model generates $k$ different entries in a single response (Zhang et al., 2024a).
  • Dynamic Few-Shot Examples: Different few-shot examples are randomly selected for each call (using a larger seed set) (Li et al., 2022).
@@
Prompting GPT-4o Llama-3.1-70B-Base
Ind. Persona Seq. In-One Dyn. Fewshot Independent
Enron 0.574 0.580 0.511 0.363 0.511 0.350
GSM8K 0.427 0.308 0.398 0.347 0.463 0.313
Table 2 shows that base models generally yield higher diversity than almost all prompting methods in the two domains, as measured by the average embedding distance. The exception is persona prompting on GSM8K — though this diversity arises more from flavor text differences due to personas rather than actual content.
We thus find that base models are generally more diverse than instruction-tuned models and that temperature increases and prompting methods are insufficient to bridge the gap, motivating our usage of base models in the first stage of BARE.
Figure 4: Indistinguishability Rate for Llama-3.1-70B Base and Instruct data generation methods across various datasets. Llama-3.1-70B-Instruct is almost uniformly better at generating examples that appear in-domain when compared alongside real-world data.
@@
Figure 4 presents our results. In general, the instruction-tuned model has a higher indistinguishability rate (IR) (see Section 3.1), indicating that it is better at producing generations that resemble high-quality data. Some IRs are well above 75%, which is not unexpected: if a model consistently generates data that aligns with the most common patterns in the real-world distribution, it becomes difficult to distinguish from actual data. Since the real-world entries often also include non-modal (less frequent) samples, a discriminator tasked with identifying lower quality data may instead misclassify these less common real-world samples as synthetic.
As mentioned above, on LCB TOP the base model repeats phrases from examples in the prompt, leading to a higher IR as the repeated examples are real. Consequently, while individual base generations are technically more realistic, their shortcomings are captured in the diversity metric.
We thus find that the superior instruction following capabilities of instruction-tuned models generates more realistic data, motivating our usage of instruction-tuned models in the second stage of BARE.
@@ 4 Base-Refine (BARE)
Figure 5: BARE combines base models with instruction tuned models. Instruction-tuned models provide high-quality but low-diversity data, while base models provide low-quality but high-diversity data. With minimal seed examples, BARE independently generates a diverse initial set of data points with a base model and refines each entry individually with an instruction-tuned model to create a high-quality, high-diversity dataset. In this example of a real grade school math problem generation, the Llama-3.1-70B-Base model hallucinates in its answer to its own question. The refiner (Llama-3.1-70B-Instruct) recognizes this and disambiguates the question and corrects the reasoning.
Building on the relative advantages of base and instruction tuned models we introduce BARE (Figure 5). BARE is a practical few-shot synthetic data generation method that combines the diversity of base models with the fidelity of instruction tuned models. BARE uses a base model to generate an initial set of diverse but potentially lower quality data from very little seed data. An instruction-tuned model then individually refines each example from the initial set, improving it according to specific criteria (e.g., realism, correctness) while keeping the original concept. The refinement retains the overall diversity of the initial set while exerting greater control over the quality of the final generations.
Importantly, diversity is elicited from the inherent properties of the base model rather than a large initial seed set, allowing for greatly improved data efficiency. Seed examples are only necessary to ensure the base model follows formatting (though they can also be included in the refine step). In our experiments, we use just three few-shot examples. In addition, we intentionally use very general prompts for BARE to demonstrate its flexibility, underscoring the potential for even greater improvement with tailored prompts. Representative prompts are included in Appendix D.
@@ 5 Evaluation
We evaluate BARE on diversity, data quality, and downstream utility across the same domains and baselines presented in Section 3. BARE uses Llama-3.1-70B-Base for generation and Llama-3.1-70B-Instruct for refinement. We also experiment with the Llama 3.1 8B family and GPT-4o as the refiner (Dubey et al., 2024; OpenAI, 2024c). Sampling details are in Appendix A.
To evaluate the downstream utility of BARE, we LoRA (Hu et al., 2021) fine-tune Llama-3.1-8B-Instruct for 4 epochs using the generated data, except for GSM8K where Llama-3.2-1B-Instruct (Dubey et al., 2024) is fine-tuned instead due to high baseline performance of the 8B model. Other fine-tuning hyperparameters are in Section B.1.1. The fine-tuned models are evaluated on a static $n=100$ test set for HotpotQA and PubMedQA, a static $n=500$ test set for GSM8K, and the full $n=442$ test set for LCB TOP. In this section, we will focus only on the generative tasks; the downstream evaluation process for classification tasks is presented in Section B.1.2 and detailed results for all domains can be found in Section B.2.
Figure 6: Distribution of pairwise embedding cosine similarity scores for Llama-3.1-70B-Base and Llama 3.1 70B BARE generations for GSM8K (top) and LCB TOP (bottom). The distributions are extremely similar for both tasks, indicating that refinement retains the diversity of base generations.
Figure 7: Indistinguishability Rate for data generated by Llama-3.1-70B-Base and BARE across various datasets. BARE consistently improves the quality of data generated from base models.
@@
We begin by comparing the quality/diversity trends of BARE with other methods. From the histograms in Figure 7, we can see that BARE effectively does not change the similarity distribution of generated data when compared to the base model at all. Detailed results with average embedding similarity scores can be found in Section B.2.
At the same time, we see from Figure 7 that BARE leads to a monotonic increase in the IR for every domain - suggesting that it is able to lift the quality of the generations to be on par or in some cases even surpass directly sampling from an instruct model. Combined, the IR and diversity measures indicate that BARE is capable of leveraging the diversity of base models and quality of instruction-tuned models in its end generations.
@@
We now show the utility of BARE-generated datasets as a whole. In Figure 8, we demonstrate the accuracy of a model fine-tuned on datasets generated using different methods. BARE-generated data leads to better downstream models than data from either Base or Instruct models almost uniformly, and across all domains training with BARE-generated data leads to the highest model accuracy.
Surprisingly, in Figure 9, we find that BARE using the small Llama-3.1-8B base model to generate data and GPT-4o as a refiner, outperforms all advanced prompting methods discussed in Section 3.3 applied to GPT-4o directly. These clear results showcase BARE’s ability to out-perform existing methods for high-quality, diverse data generation in few-shot settings.
Figure 8: Downstream accuracy of BARE compared to Llama-3.1 Base and Instruct models. Accuracy is measured on Llama-3.1-8B-Instruct Model (HotpotQA, PubMedQA, LCB TOP) and Llama-3.2-1B-Instruct Model (GSM8K) finetuned on synthetic data generated using Base, Instruct, and BARE methods. Note that GSM8K 8B Instruct results are not shown as generations derailed. In general, we find that BARE out-performs directly using the base and instruction fine-tuned models.
Figure 9: Downstream accuracy for BARE versus other diversification methods. Fine-tuned model accuracy on HotpotQA, PubMedQA, GSM8K, and LCB TOP for data generated from prompting methods using GPT-4o. Prompting methods have mixed effects on downstream performance when compared against standard sampling. Alternatively, BARE using just with the Llama-3.1-8B Base as the generator and GPT-4o as the refiner significantly outperforms all over methods using GPT-4o.
On RAFT domains, BARE improves upon the standard state-of-the-art pipeline for fine-tuning LLMs for RAG (RAFT) by up to 18.4%, as seen with the 8B family on HotpotQA (standard RAFT is represented in our Instruct generation results). BARE with both model families also outperforms existing RAFT pipelines on PubMedQA.
On GSM8K, BARE is the only method that provides useful training data. The un-trained model performance was 21.8%, and fine-tuning on BARE-generated data achieves accuracies of 24.9% and 32.8% with the Llama 8B and 70B families, respectively. Accuracy when training with data generated by single model methods either decreased accuracy or had little difference. In fact, training on data generated by Llama-3.1-70B-Instruct led to an accuracy of just 17.8%, which BARE with Llama-3.1-70B-Base refined by GPT-4o outperforms by over 2×2$\times$2 × (35.8% accuracy).
@@
We note that with the 70B family on both HotpotQA and LCB, BARE-generated data leads to less of an improvement than the instruct-generated data. In these cases, using GPT-4o as a refiner instead of Llama-3.1-70B-Instruct does lead to stronger results, indicating that the choice of refiner can strongly influence data quality: the Llama model was unable to sufficiently refine the base model generations, but the stronger GPT-4o was able to do so. We show detailed results of refining with GPT-4o in all domains in Section B.2, emphasizing BARE’s consistently strong data generation capabilities. For more detailed discussion, see Appendix C.
Comparison to Real-World Data.
To further show the quality of our synthetic data, we took the real GSM8K training data (the only domain of these 4 which has a real train set) and randomly sampled 1,000 examples from which to fine-tune the Llama-3.2-1B-Instruct model. The resulting trained model had a performance of $26.6\%$, surpassing the 8B/70B Base and Instruct models individually. Interestingly, though, it was surpassed by almost all BARE methods, most prominently when GPT-4o was used as a refiner with Llama-3.1-70B-Base (35.8%). While one might expect the real training data should be better than a synthetically generated dataset based on a small sample, the quality of synthetically generated data with an emphasis on quality and diversity can lead to better generalization then the original human generated training data. This is in line with other works’ findings for synthetic sets generated from large samples (Liu et al., 2023).
Temperature & Scale Ablations. We also find that temperature ablations do not meaningfully affect the utility of instruction-tuned model generations, especially when compared to gains using BARE; for details, see Section B.3. Additionally, we found that BARE’s strong performance holds when scaling the amount of generated data; for details, see Section B.5.
Figure 10: Accuracy at varying seed (few-shot) sizes of a Llama-3.2-1B-Instruct Model finetuned on synthetic data generated using Instruct and BARE methods on GSM8K. BARE performance is consistent throughout while Instruct improves with seed set size before converging with BARE.
@@
To verify BARE’s performance comes from the use of base models rather than the multi-step setup, we replace the base model in the first step of BARE with an instruction-tuned model on the GSM8K task. Switching Llama-3.1-70B-Base to Instruct drops the accuracy on the test set from $29.8\%$ to $25.4\%$ when refining with Llama-3.1-70B-Instruct and from $35.8\%$ to $30.8\%$ with GPT-4o. At the same time, the average pairwise similarity before and after refinement remains similar, indicating that diversity must be introduced in the first stage. Having established that base models are most effective at seed efficient diversity, we conclude that the use of base models is a key reason for the utility of BARE. Detailed results from this ablation are in Section B.4.
Seed Data Scale Ablation.
We additionally investigate the advantage of BARE compared to instruct-only methods as more seed data is made available. We evaluate results with BARE and Instruct using the Llama 3.1 70B family on GSM8K, ranging the seed set size from $3$ to $3000$ and continuing to generate 1,000 examples (Figure 10). Each generation randomly selects three examples from the seed set for the prompt.
BARE consistently outperforms Instruct generation for small seed sets, highlighting the potential of BARE in making high-quality synthetic data generation easier to adopt and realistic in low-data domains. The methods converge and plateau for large seed set sizes, indicating that diversity sourced from the base model’s priors in BARE is always at least as effective as diversity from curated seed sets. In our experiments, convergence between BARE and Instruct generation occurred at around $100$ examples, or $10\%$ of the generated dataset size. We leave as future work the validation of the convergence point at larger dataset scales.
@@ 6 Conclusion
In this work we investigated the potential of using base models for few-shot synthetic data generation. Through this investigation, we find that base models are an effective source of diversity even when given minimal seed data, making them an attractive choice for data efficiency. These insights motivate the design of a system novelly leveraging base models in synthetic data generation, BARE.
Through extensive experiments, we validate the importance of each step in BARE and demonstrate its ability to preserve base model diversity while enhancing output quality. The use and analysis of base models specifically in BARE opens an avenue for much future work as almost no prior work to our knowledge has studied their innate value (before post-training).
Moreover, by fine-tuning on BARE-generated data for various domains, we underscore BARE’s practical utility, consistently outperforming existing synthetic data generation methods on downstream tasks such as GSM8K and LiveCodeBench, in addition to RAFT, for which we set a new SOTA.
Acknowledgments and Disclosure of Funding
@@
Sky Computing Lab is supported by gifts from Accenture, AMD, Anyscale, Cisco, Google, IBM, Intel, Intesa Sanpaolo, Lambda, Microsoft, NVIDIA, Samsung SDS, SAP, and VMware. We would additionally like to acknowledge Databricks for their generous compute support.
@@
Appendix A Sampling Details
During data generation, all base models were sampled at a default temperature of $0.7$. For instruction-tuned-only generations we sampled at the highest temperature at which we experimentally found data generation to still be coherent, which is $1.0$ for Llama models and $1.2$ for GPT-4o. However, for Enron, we sample from GPT-4o at a temperature of $1.0$ to maintain generation coherence. Similarly, for Enron and Newsgroups, we sample for Llama-3.1-8B-Instruct at a temperature of $0.7$. Instruction-tuned models used during refinement are always sampled at a temperauter of $0.7$.
Appendix B Additional Results
B.1 Downstream Evaluation - Additional Details
B.1.1 Fine-Tuning Task Hyperparameters